London purchased a piece of real estate last year for $81,200. The real estate is now worth $100,000. If London needs to have a total return of 0.24 during the year, then what is the dollar amount of income that she needed to have to reach her objective? Round to two decimal places.

0 0

> I think if you are indeed interested in such special house I would say you just call the good listing agent whose name is on the sign/ad that you see on good real estate websites - they will provide you with relevant information on the house and give you a property tour. Referrals and expert reviews available on such sites also help you in making the right decision.

Anthony Constantinou Wed, Dec 8, 2021 8:31 AM

Add a comment Improve this question Transcribed image text
Answer #1

Total return=(End value-Beginning value+Income)/Beginning value  

0.24=(100,000-81200+Income)/81200

(0.24*81200)=18800+Income

Income=19488-18800

=$688